# Small Neutrino Masses from a Decoupled Singlet Scalar Field

### v1

@@ Abstract
Dark matter direct detection: When the lighter neutral scalar from the inert doublet is dark matter, we can avoid the strong bound from direct detection if the mass splitting between $A_{0}$ and $H_{0}$ is large enough, i.e. $|\Delta|\gtrsim 100\,{\rm keV}$. Then, the cross section for spin-independent scattering between dark matter and nucleus is dominated by the tree-level Higgs exchange Barbieri:2006dq ; LopezHonorez:2006gr , as follows,
$\sigma^{\rm SI}_{A_{0}-N}=\frac{(\lambda_{L}-\lambda_{5,\,{\rm eff}})^{2}\mu^{2}_{N}}{4\pi m^{2}_{A_{0}}m^{4}_{h}A^{2}}(Zf_{p}+(A-Z)f_{n})^{2}$ (10)
@@
where $Z,A-Z$ are the numbers of protons and neutrons in the detector nucleus, $\mu_{N}$ is the reduced mass for the DM-nucleus system, and $f_{p,n}$ are the nucleon form factors Hoferichter:2015dsa ; Junnarkar:2013ac ; Kim:2023pwf .
@@
The DM relic density is determined by the DM annihilations into a pair of SM fermions, $WW,ZZ$ and $h_{\rm SM}h_{\rm SM}$ with $h_{\rm SM}$ being the SM Higgs boson Barbieri:2006dq ; LopezHonorez:2006gr ; Gong:2012ri . In particular, there is a room to adjust $\lambda_{L}$ to explain the DM relic density for relatively heavy DM due to $A_{0}A_{0}\to h_{\rm SM}h_{\rm SM}$ while the direct detection bound is satisfied Belyaev:2016lok . DM masses above about $100\,{\rm GeV}$ are consistent with the bounds from direct detection, as far as $|\lambda_{L}-\lambda_{5,\,{\rm eff}}|\lesssim 10^{-2}$, apart from the resonance. But, if $|\lambda_{L}-\lambda_{5,\,{\rm eff}}|\lesssim 10^{-3}$, as required near resonances, the one-loop corrections to the DM-nucleus cross section are relevant Abe:2015rja .
@@
Thermal leptogenesis: Baryogenesis can be realized in our model via thermal leptogenesis Fukugita:1986hr , due to the decays of RH neutrinos with hierarchical masses. As the asymmetries generated from the decays of the heavier RH neutrinos, $N_{2,3}$, are negligible due to strong washout effects by scattering, we focus on the $N_{1}$ contribution to the lepton asymmetry. The baryon-to-photon ratio converted from the final $B-L$ asymmetry is given by
@@
Here, $\Gamma_{1}$ is the decay rate for $N_{1}$, given by $\Gamma_{1}\simeq\frac{M_{N,1}}{8\pi}(y^{\dagger}_{N}y_{N})_{11}$ for $m_{0}\ll M_{N,1}$, and $H(z_{1})$ is the Hubble parameter. In the Casas-Ibarra parametrization for the Yukawa matrix $y_{N}$ Casas:2001sr , we can express $(y^{\dagger}_{N}y_{N})_{ij}=\sqrt{\Lambda_{i}\Lambda_{j}}{\widetilde{m}}_{ij}$ in terms of the effective neutrino mass matrix ${\widetilde{m}}\equiv RD_{{\cal M}_{\nu}}R^{\dagger}$ with $R$ being arbitrary complex rotation matrix, $D_{{\cal M}_{\nu}}={\rm diag}(m_{\nu_{1}},m_{\nu_{2}},m_{\nu_{3}})$ and $\Lambda_{i}=\frac{4\pi^{2}\xi_{i}M_{N,i}}{\lambda_{5,{\rm eff}}v^{2}_{H}}$ Hugle:2018qbw . Here, $\xi_{i}=\big{(}\frac{1}{8}\frac{M^{2}_{N,i}}{m^{2}_{H_{0}}-m^{2}_{A_{0}}}[L(m^{2}_{H_{0}})-L(m^{2}_{A_{0}})]\big{)}^{-1}$ with $L(m^{2})=\frac{m^{2}}{m^{2}-M^{2}_{N,i}}\ln(m^{2}/M^{2}_{N,i})$ are order one parameters in most of the parameter space. Thus, we get the decay parameter as $K_{1}\simeq\frac{2\pi^{2}\xi_{1}}{\lambda_{5\,{\rm eff}}}\sqrt{\frac{45}{64\pi^{5}g_{*}}}\frac{M_{P}}{v^{2}_{H}}{\widetilde{m}}_{11}$ for $m_{0}\ll M_{N,1}$.
@@
Figure 4: Parameter space for RH neutrino mass $M_{N}$ vs $\lambda_{5,{\rm eff}}$, showing $\eta_{B}>6.1\times 10^{-10}$ in purple region.
@@
Conclusions: We presented a unified solution with $Z_{4}$ discrete Peccei-Quinn symmetry for small neutrino masses and stability of dark matter in the radiative seesaw scenarios. Once the $Z_{4}$ symmetry is broken to $Z_{2}$ by the VEV of an extra singlet scalar field $\varphi$ much below the mass scale of the singlet scalar $S$, a small mixing between the dark sector neutral scalars is generated. Then, we showed that small neutrino masses are obtained from one-loop diagrams with RH neutrinos and inert doublet scalars with small mass splittings and identified the parameter space where observed neutrino masses are explained and the lighter scalar of the inert doublet becomes a dark matter candidate. From the measurement of the absolute neutrino masses at the level of $0.01\,{\rm eV}$, direct detection searches for the inert doublet dark matter and the successful leptogenesis, we showed that a new mass scale of the singlet scalar $S$ is expected to be up to $\sim 10^{16}\,{\rm GeV}$.
@@ Acknowledgments
The work is supported in part by Basic Science Research Program through the National Research Foundation of Korea (NRF) funded by the Ministry of Education, Science and Technology (NRF-2022R1A2C2003567 and RS-2024-00341419 (JKK)).

### v2

@@ V Dark matter direct detection
When the lighter neutral scalar from the inert doublet is dark matter, we can avoid the strong bound from direct detection if the mass splitting between $A_{0}$ and $H_{0}$ is large enough, i.e. $|\Delta|\gtrsim 100\,{\rm keV}$. Then, the cross section for spin-independent scattering between dark matter and nucleus is dominated by the tree-level exchanges of Higgs-like scalars Barbieri:2006dq ; LopezHonorez:2006gr , as follows,
$\sigma^{\rm SI}_{A_{0}}=\frac{\lambda_{L5}^{2}\mu^{2}_{N}(Zf_{p}+(A-Z)f_{n})^{2}}{4\pi m^{2}_{A_{0}}A^{2}}\bigg{(}\frac{c^{2}_{\alpha}}{m^{2}_{h_{1}}}-\frac{s^{2}_{\alpha}}{m^{2}_{h_{2}}}\bigg{)}^{2}$ (10)
@@
where $\lambda_{L5}\equiv\lambda_{L}-\lambda_{5,\,{\rm eff}}$, $m_{h_{1,2}}$ are the masses for SM Higgs-like and dark Higgs-like scalars, $Z,A-Z$ are the numbers of protons and neutrons in the detector nucleus, $\mu_{N}$ is the reduced mass for the DM-nucleus system, and $f_{p,n}$ are the nucleon form factors Hoferichter:2015dsa ; Junnarkar:2013ac ; Kim:2023pwf .
@@
The DM relic density is determined by the DM annihilations into a pair of SM fermions, $WW,ZZ$ and $h_{i}h_{j}$ with $i,j=1,2$ Barbieri:2006dq ; LopezHonorez:2006gr ; Gong:2012ri . If $H_{0},H^{\pm}$ have comparable masses to $m_{A_{0}}$, we also need to include the co-annihilation channels associated with them as eq. (57) in the appendix. In particular, there is room to adjust $\lambda_{L}$ and dark Higgs-portal couplings, denoted by ${\cal L}_{\rm int}\supset-\frac{1}{2}\sum_{i=1,2}\lambda_{H_{i}\varphi}|H_{i}|^{2}\varphi^{2}$, to explain the DM relic density for relatively heavy DM due to $A_{0}A_{0}\to h_{i}h_{j}$ while the direct detection bound is satisfied Belyaev:2016lok . DM masses above about $100\,{\rm GeV}$ are consistent with the bounds from direct detection, as far as $|\lambda_{L5}(c^{2}_{\alpha}-s^{2}_{\alpha}m^{2}_{h_{1}}/m^{2}_{h_{2}})|\lesssim 10^{-2}$, apart from the resonance, and the bounds from the SM Higgs decays can be easily satisfied for $m_{A_{0}},m_{h_{2}}\gtrsim m_{h_{1}}/2=62.5\,{\rm GeV}$. But, for a smaller value of $|\lambda_{L5}|$, as required near resonances, the one-loop corrections to the DM-nucleus cross section are relevant Abe:2015rja .
In Fig. 5, we depict the DM relic density as a function of DM mass for $\lambda_{H_{2}\varphi}=0,0.01,0.03$, in red solid, blue dashed and dotted lines, respectively. We chose a zero Higgs mixing, $\alpha=0$, the singlet parameters are $\lambda_{H_{1}\varphi}=0$, $v_{\varphi}=10^{4}\,{\rm GeV}$ and $m_{h_{2}}=m_{\rho}=80\,{\rm GeV}$, and the parameters for the inert doublet are $\lambda_{L5}=0.01$, $\Delta=10\,{\rm MeV}$ and $m_{H^{\pm}}=m_{0}+1\,{\rm GeV}$. Thus, as $\lambda_{H_{2}\varphi}$ increases, the correct relic density can be explained for a heavier DM mass far beyond $100\,{\rm GeV}$. The results are insensitive to $\Delta$ as far as $\Delta\lesssim T_{f}$ where $T_{f}$ is the freeze-out temperature given by $T_{f}\simeq m_{0}/20$. Moreover, we remark that a nonzero $\lambda_{H_{1}\varphi}$ also contributes to the dark matter annihilation into $\varphi\varphi$ with the SM Higgs exchange, but its contribution is suppressed for a small $|\lambda_{L5}|$.
Figure 5: DM relic density as a function of DM mass $m_{0}$, depending on the inert doublet portal coupling, $\lambda_{H_{2}\varphi}$.
VI Thermal leptogenesis
@@
Baryogenesis can be realized in our model via thermal leptogenesis Fukugita:1986hr , due to the decays of RH neutrinos with hierarchical masses. The RH neutrinos can be once in thermal equilibrium with the SM plasma in the early universe due to their Yukawa couplings to the inert doublet or the spurion scalar, which mixes with the SM Higgs. As the asymmetries generated from the decays of the heavier RH neutrinos, $N_{2,3}$, are negligible due to strong washout effects by scattering, we focus on the $N_{1}$ contribution to the lepton asymmetry. The baryon-to-photon ratio converted from the final $B-L$ asymmetry is given by
@@
Here, $\Gamma_{1}$ is the decay rate for $N_{1}$, given by $\Gamma_{1}\simeq\frac{M_{N,1}}{8\pi}(y^{\dagger}_{N}y_{N})_{11}$ for $m_{0}\ll M_{N,1}$, and $H(z_{1})$ is the Hubble parameter. In the Casas-Ibarra parametrization for the Yukawa matrix $y_{N}$ Casas:2001sr , we can express $(y^{\dagger}_{N}y_{N})_{ij}=\sqrt{\Lambda_{i}\Lambda_{j}}{\widetilde{m}}_{ij}$ in terms of the effective neutrino mass matrix ${\widetilde{m}}\equiv RD_{{\cal M}_{\nu}}R^{\dagger}$ with $R$ being arbitrary complex rotation matrix, $D_{{\cal M}_{\nu}}={\rm diag}(m_{\nu_{1}},m_{\nu_{2}},m_{\nu_{3}})$, and $\Lambda_{i}=\frac{4\pi^{2}\xi_{i}M_{N,i}}{\lambda_{5,{\rm eff}}v^{2}_{H}}$ Hugle:2018qbw . Here, $\xi_{i}=\big{(}\frac{1}{8}\frac{M^{2}_{N,i}}{m^{2}_{H_{0}}-m^{2}_{A_{0}}}[L(m^{2}_{H_{0}})-L(m^{2}_{A_{0}})]\big{)}^{-1}$ with $L(m^{2})=\frac{m^{2}}{m^{2}-M^{2}_{N,i}}\ln(m^{2}/M^{2}_{N,i})$ are order one parameters in most of the parameter space. Thus, we get the decay parameter as $K_{1}\simeq\frac{2\pi^{2}\xi_{1}}{\lambda_{5\,{\rm eff}}}\sqrt{\frac{45}{64\pi^{5}g_{*}}}\frac{M_{P}}{v^{2}_{H}}{\widetilde{m}}_{11}$ for $m_{0}\ll M_{N,1}$.
Figure 6: Parameter space for RH neutrino mass $M_{N}$ vs $\lambda_{5,{\rm eff}}$, showing $\eta_{B}>6.1\times 10^{-10}$ in purple region. The contour lines show the parameter space for the heaviest neutrino mass equal to $m_{\nu}=0.05\,{\rm eV}$.
@@
VII Conclusions
We presented a unified solution with $Z_{4}$ discrete symmetry for small neutrino masses and stability of dark matter in the radiative seesaw scenarios. Once the $Z_{4}$ symmetry is broken to $Z_{2}$ by the VEV of the spurion singlet scalar $\varphi$ much below the mass of the singlet scalar $S$, a small effective lepton number violating interaction between two Higgs doublets is generated, being responsible for a small mass splitting of the neutral inert scalars at tree level and small neutrino masses at one-loop level. We showed that the spurion scalar opens up a new parameter space for the correct relic density for dark matter and a successful leptogenesis is achieved.
@@ Acknowledgments
The work is supported in part by Basic Science Research Program through the National Research Foundation of Korea (NRF) funded by the Ministry of Education, Science and Technology (NRF-2022R1A2C2003567 and RS-2024-00341419 (JKK)). This research was supported by the Chung-Ang University research grant in 2019.